OpenStudy (anonymous):

On a coordinate plane, the coordinates of vertices R and T for a polygon are R(−6, 2) and T(5, 2). What is the length of Side RT of the polygon? I need help on this im bad at area. 4 6 7 11

OpenStudy (anonymous):

You simply calculate the distance between coordinates R and T, like this: \[r = \sqrt{(-6-5)^2 + (2-2)^2} = \sqrt{11^2} = 11\]

OpenStudy (anonymous):

i knew it was 11

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

anytime ;)